# Deformation quantization and perverse sheaves

### v1

@@ 5.2. Aside: twisted de Rham complex
By [SS14, Proposition 2] the induced morphism
@@
$(\mathrm{DR}_{f},\hbar\partial_{\hbar}-\hbar^{-1}f)\cong\widehat{\mathrm{RH}}^{-1}(\phi_{f}).$
@@
Such an isomorphism was established in [Sab10, SS14] on the level of cohomology sheaves and was recently upgraded in [Sch23] to an isomorphism of differential perverse sheaves, by different methods than above.
@@ 5.3. Thom–Sebastiani theorem
In this section, for any complex manifold $X$, we will consider ^$\widehat{\mathcal{E}}_{X\times\mathbf{C}}$ as a sheaf on $\mathrm{J}^{1}X\subseteq\mathrm{P}^{\ast}(X\times\mathbf{C})$. We denote by $\alpha\colon\mathrm{J}^{1}X_{1}\times\mathrm{J}^{1}X_{2}\to\mathrm{J}^{1}(X_{1}\times X_{2})$ the addition map. Consider the corresponding transfer $(\alpha^{-1}\widehat{\mathcal{E}}_{X_{1}\times X_{2}\times\mathbf{C}_{t}},\widehat{\mathcal{E}}_{X_{1}\times\mathbf{C}_{t_{1}}\times X_{2}\times\mathbf{C}_{t_{2}}})$-bimodule (see [SKK73, Section 1.3] for the general notion of transfer bimodules)
@@
Indeed, for the first map this is shown in [MSS20, Section 1.1] and for the composite this is (2).
The $V$-filtrations on ~$\widetilde{\mathcal{B}}_{f_{i}}$ induce by convolution a filtration on ~$\widetilde{\mathcal{B}}_{f_{1}}\stackrel{{\scriptstyle\mathcal{D}}}{{\boxtimes}}\widetilde{\mathcal{B}}_{f_{2}}$. It is shown in [MSS20, Theorem 1.2] and [SS, Lemma 11.8.2] that the natural map
@@
Unwinding the construction of $M_{f}$ in 5.7 via microlocalization, the top line of the diagram under consideration may be computed as the image under $\widehat{\mathrm{RH}}^{-1}$ of a sequence of morphisms of perverse sheaves with automorphism:
@@ 6.1. Exact two-forms on complex-analytic spaces
We will use the notions related to d-critical loci from [Bra+15, Joy15] which we now briefly recall. Let us first explain how to construct exact two-forms of degree $-1$ on critical loci.
@@
We have the following important theorems proved in [Joy15]:
  • Given an embedding $\Xi\colon(R,U,f,i)\hookrightarrow(S,V,g,j)$ of critical charts there is a canonical quadratic form $q_{\Xi}$ on the normal bundle $N_{\Xi}|_{R}$ of $U\subset V$. The quadratic vector bundle $N_{\Xi}|_{R}$ defines a $\mathbf{Z}/2$-torsor $P_{\Xi}$ on $R$ of orientations of $q_{\Xi}$. See [Joy15, Definition 2.26].
  • An embedding of critical charts $(R,U,f,i)\hookrightarrow(S,V,g,j)$ is, locally, of the form $(R,U,f,i)\hookrightarrow(R,U\times\mathbf{C}^{n},f\boxplus(z_{1}^{2}+\dots+z_{n}^{2}),j)$, where $U\hookrightarrow U\times\mathbf{C}^{n}$ is given by $u\mapsto(u,0)$. See [Joy15, Proposition 2.22].
  • Given critical charts $(R,U,f,i)$ and $(S,V,g,j)$ and an intersection point $x\in R\cap S$ we can find near $x$ another critical chart $(T,W,h,k)$ and local embeddings $(R,U,f,i)\hookrightarrow(T,W,h,k)$ and $(S,V,g,j)\hookrightarrow(T,W,h,k)$. See [Joy15, Theorem 2.20].
@@ 6.3. Lagrangian intersections
In this section we explain the main example of a d-critical locus in this paper given by a Lagrangian intersection. Let $S$ be a holomorphic symplectic manifold and $L,M\subset S$ two Lagrangian submanifolds. Consider the intersection $L\cap M$ as a complex analytic space. The following is a generalization of 6.3 proved in [Bus14].
@@
Moreover, it is clear that this d-critical structure on $L\cap M$ coincides with the one constructed in [Bus14, Theorem 3.1], where the virtual canonical bundle is identified in the same way. ∎
@@ 6.4. DT sheaf
Let $(X,s)$ be an oriented d-critical locus. The paper [Bra+15] constructs a canonical perverse sheaf $\phi_{X}\in\mathrm{Perv}_{\mathrm{aut}}(X)$ equipped with an automorphism on $X$, which we call the DT sheaf, locally modeled on the sheaf of vanishing cycles. Let us briefly recall its construction.
@@
Let $\Xi\colon(R,U,f,i)\hookrightarrow(S,V,g,j)$ be an embedding of critical charts. Then [Bra+15, Theorem 5.4] shows that there is the stabilization isomorphism
@@
  • Locally near a point of $R$ identify $V\cong R\times W$ and $g\mapsto f\boxplus q$ for a quadratic form $q\colon W\rightarrow\mathbf{C}$ using [Joy15, Proposition 2.22], so that under this isomorphism the embedding becomes $U\hookrightarrow U\times W$ given by $u\mapsto(u,0)$. In this local model the isomorphism $\Theta_{\Xi}$ is equal to the composite
@@
  • By [Joy15, Theorem 2.20] any two critical charts $(R,U,f,i)$, $(S,V,g,j)$ are locally related by a zig-zag $(R,U,f,i)\hookrightarrow(T,W,h,k)\hookleftarrow(S,V,g,j)$ of critical embeddings. The isomorphism of the local models for $\phi_{X}$ in charts $(R,U,f,i)$ and $(S,V,g,j)$ is given by applying the stabilization isomorphism (12) to the two critical embeddings.
@@ 7. Comparison of the two perverse sheaves
Throughout this section $S$ denotes a complex symplectic manifold, $\rho\colon Y\rightarrow S$ its contactification, $L,M\subset S$ Lagrangian submanifolds equipped with orientation data and $\Lambda_{L},\Lambda_{M}\subset Y$ their Legendrian lifts. The Lagrangian intersection $L\cap M$ has a structure of an oriented d-critical locus by 6.7. The goal of this section is to prove the following result.
@@
$\mathsf{M}^{\lambda}_{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathsf{M}^{\mu}_{M}\cong\mathrm{RH}^{-1}(\phi_{L\cap M}\otimes\mathbf{C}_{\mu+\lambda}),$
@@
Let $\dim L=\dim M=n$.
@@
The manifold $S$ admits local polarizations. We use them to get a local description of $(Y\rightarrow S,\Lambda_{L}\rightarrow L,\Lambda_{M}\rightarrow M)$ following the terminology from [Bus14, Section 2]:
@@ 7.2. Local description of DQ modules
  • In the $L$-chart we have $\mathsf{E}_{Y}=\widehat{\mathcal{E}}^{\sqrt{v}}_{L\times\mathbf{C}}$ and $\mathsf{W}_{S}=\widehat{\mathcal{W}}^{\sqrt{v}}_{L}$. The modules are
@@
  • In the $M$-chart we have $\mathsf{E}_{Y}=\widehat{\mathcal{E}}^{\sqrt{v}}_{M\times\mathbf{C}}$ and $\mathsf{W}_{S}=\widehat{\mathcal{W}}^{\sqrt{v}}_{M}$. The modules are
@@
We use the gluing data for the perverse sheaf $\phi_{L\cap M}$ as described in [Bus14, Section 2.3]. Cover $S$ by open sets which admit polarizations and consider them as $L$-charts. Given an intersection of two such $L$-charts $U_{1},U_{2}$, we may cover $U_{1}\cap U_{2}$ by open sets $\{V_{a}\}$ with polarizations such that for every $a$ the polarization on $V_{a}$ is transverse both to the polarization on $U_{1}$ and on $U_{2}$. We regard $V_{a}$ as $M$-charts.
@@
It is enough to check commutativity of the diagram locally. Consider a point $x\in L\cap M$. We have that $\Xi\colon(L\cap M,L,f,i)\rightarrow(L\cap M,L\times M,h,i)$ is a critical embedding. Therefore, by [Joy15, Proposition 2.22] in a neighborhood of $x$ there is a local isomorphism $L\times M\cong L\times W$, where $W=\mathrm{T}_{x}M$ is equipped with the Maslov quadratic form $q=q(\mathrm{T}_{M},\mathrm{T}_{\pi_{1}},\mathrm{T}_{\pi_{2}})\colon W\rightarrow\mathbf{C}$, such that under this isomorphism $h\mapsto f\boxplus q$ and the embedding $\Xi\colon L\rightarrow L\times M$ becomes $(\mathrm{id}\times 0)\colon L\rightarrow L\times W$. The $\mathbf{Z}/2$-torsor $P_{\Xi}$ in a neighborhood of $x$ canonically identifies with $P_{q}$.
@@ 7.4. Example: a clean Lagrangian intersection
described in [Joy15, Theorem 2.28 (i)] for a general d-critical locus. By [Bus14, Theorem 3.1] we may additionally identify

### v2

@@ 5.2. Aside: twisted de Rham complex
By [SabbahSaito, Proposition 2] the induced morphism
@@
Such an isomorphism was established in [Sabbah, SabbahSaito] on the level of cohomology sheaves and was recently upgraded in [Schefers] to an isomorphism of differential perverse sheaves, by different methods than above.
@@ 5.3. Thom–Sebastiani theorem
In this section, for any complex manifold $X$, we will consider ^$\widehat{\mathcal{E}}_{X\times\mathbf{C}}$ as a sheaf on $\mathrm{J}^{1}X\subseteq\mathrm{P}^{\ast}(X\times\mathbf{C})$. We denote by $\alpha\colon\mathrm{J}^{1}X_{1}\times\mathrm{J}^{1}X_{2}\to\mathrm{J}^{1}(X_{1}\times X_{2})$ the addition map. Consider the corresponding transfer $(\alpha^{-1}\widehat{\mathcal{E}}_{X_{1}\times X_{2}\times\mathbf{C}_{t}},\widehat{\mathcal{E}}_{X_{1}\times\mathbf{C}_{t_{1}}\times X_{2}\times\mathbf{C}_{t_{2}}})$-bimodule (see [SatoKawaiKashiwara, Section 1.3] for the general notion of transfer bimodules)
@@
Indeed, for the first map this is shown in [MaximSaitoSchurmann, Section 1.1] and for the composite this is (2).
The $V$-filtrations on ~$\widetilde{\mathcal{B}}_{f_{i}}$ induce by convolution a filtration on ~$\widetilde{\mathcal{B}}_{f_{1}}\stackrel{{\scriptstyle\mathcal{D}}}{{\boxtimes}}\widetilde{\mathcal{B}}_{f_{2}}$. It is shown in [MaximSaitoSchurmann, Theorem 1.2] and [SabbahSchnell, Lemma 11.8.2] that the natural map
@@
Unwinding the construction of $M_{f}$ in 5.7 via microlocalization, the top line of the diagram under consideration may be computed as the image under $\mathrm{RH}^{-1}$ of a sequence of morphisms of perverse sheaves with automorphism:
@@ 6.1. Exact two-forms on complex-analytic spaces
We will use the notions related to d-critical loci from [BBDJS, JoycedCrit] which we now briefly recall. Let us first explain how to construct exact two-forms of degree $-1$ on critical loci.
@@
We have the following important theorems proved in [JoycedCrit]:
  • Given an embedding $\Xi\colon(R,U,f,i)\hookrightarrow(S,V,g,j)$ of critical charts there is a canonical quadratic form $q_{\Xi}$ on the normal bundle $N_{\Xi}|_{R}$ of $U\subset V$. The quadratic vector bundle $N_{\Xi}|_{R}$ defines a $\mathbf{Z}/2$-torsor $P_{\Xi}$ on $R$ of orientations of $q_{\Xi}$. See [JoycedCrit, Definition 2.26].
  • An embedding of critical charts $(R,U,f,i)\hookrightarrow(S,V,g,j)$ is, locally, of the form $(R,U,f,i)\hookrightarrow(R,U\times\mathbf{C}^{n},f\boxplus(z_{1}^{2}+\dots+z_{n}^{2}),j)$, where $U\hookrightarrow U\times\mathbf{C}^{n}$ is given by $u\mapsto(u,0)$. See [JoycedCrit, Proposition 2.22].
  • Given critical charts $(R,U,f,i)$ and $(S,V,g,j)$ and an intersection point $x\in R\cap S$ we can find near $x$ another critical chart $(T,W,h,k)$ and local embeddings $(R,U,f,i)\hookrightarrow(T,W,h,k)$ and $(S,V,g,j)\hookrightarrow(T,W,h,k)$. See [JoycedCrit, Theorem 2.20].
@@ 6.3. Lagrangian intersections
In this section we explain the main example of a d-critical locus in this paper given by a Lagrangian intersection. Let $S$ be a holomorphic symplectic manifold and $L,M\subset S$ two Lagrangian submanifolds. Consider the intersection $L\cap M$ as a complex analytic space. The following is a generalization of 6.3 proved in [Bussi].
@@
Moreover, it is clear that this d-critical structure on $L\cap M$ coincides with the one constructed in [Bussi, Theorem 3.1], where the virtual canonical bundle is identified in the same way. ∎
@@ 6.4. DT sheaf
Let $(X,s)$ be an oriented d-critical locus. The paper [BBDJS] constructs a canonical perverse sheaf $\phi_{X}\in\mathrm{Perv}_{\mathrm{aut}}(X)$ equipped with an automorphism on $X$, which we call the DT sheaf, locally modeled on the sheaf of vanishing cycles. Let us briefly recall its construction.
@@
Let $\Xi\colon(R,U,f,i)\hookrightarrow(S,V,g,j)$ be an embedding of critical charts. Then [BBDJS, Theorem 5.4] shows that there is the stabilization isomorphism
@@
  • Locally near a point of $R$ identify $V\cong R\times W$ and $g\mapsto f\boxplus q$ for a quadratic form $q\colon W\rightarrow\mathbf{C}$ using [JoycedCrit, Proposition 2.22], so that under this isomorphism the embedding becomes $U\hookrightarrow U\times W$ given by $u\mapsto(u,0)$. In this local model the isomorphism $\Theta_{\Xi}$ is equal to the composite
@@
  • By [JoycedCrit, Theorem 2.20] any two critical charts $(R,U,f,i)$, $(S,V,g,j)$ are locally related by a zig-zag $(R,U,f,i)\hookrightarrow(T,W,h,k)\hookleftarrow(S,V,g,j)$ of critical embeddings. The isomorphism of the local models for $\phi_{X}$ in charts $(R,U,f,i)$ and $(S,V,g,j)$ is given by applying the stabilization isomorphism (12) to the two critical embeddings.
@@
We denote by $\phi^{\mathcal{D}}_{X}\in\mathrm{Mod}_{\mathrm{rh},\mathrm{aut}}(\mathcal{D}_{X})$ the $D$-module corresponding to $\phi_{X}$ under the Riemann–Hilbert correspondence.
@@ 7. Comparison of the two perverse sheaves
7.1. Statement
Throughout this section $S$ denotes a complex symplectic manifold, $\rho\colon Y\rightarrow S$ its contactification, $L,M\subset S$ Lagrangian submanifolds equipped with orientation data $K_{L}^{1/2},K_{M}^{1/2}$ and $\Lambda_{L},\Lambda_{M}\subset Y$ their Legendrian lifts. The Lagrangian intersection $L\cap M$ has a structure of an oriented d-critical locus by 6.7.
By 3.16 microlocalization establishes an equivalence
$\mu^{dR}_{\Lambda_{L}}\colon\mathrm{Mod}_{\Lambda,\mathrm{rh}}(\widehat{\mathsf{E}}_{Y})\xrightarrow{\sim}\gamma_{*}\mathrm{Mod}_{\mathrm{lisse}}(\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}_{L}}).$
We may canonically identify $\tilde{\Lambda}_{L}\cong\Lambda_{L}\times\mathbf{C}^{\times}\cong L\times\mathbf{C}^{\times}$. Let $i_{L}\colon L\rightarrow\tilde{\Lambda}_{L}$ be the inclusion of $1\times\mathbf{C}^{\times}$ and similarly for $i_{M}\colon M\rightarrow\tilde{\Lambda}_{M}$. As explained in Section 2.1, for a monodromic (twisted) $D$-module $M$ on $\tilde{\Lambda}_{L}$ the pullback $i_{L}^{*}M$ is a (twisted) $D$-module on $L$ equipped with an automorphism.
Fix $\lambda\in\mathbf{C}$. Then, up to isomorphism, there is a unique $\widehat{\mathsf{E}}_{Y}$-module $\mathsf{M}^{\lambda}_{L}$ simple along $\Lambda_{L}$ such that
$i_{L}^{*}\mu_{\Lambda_{L}}^{dR}(\mathsf{M}^{\lambda}_{L})\cong K_{L}^{1/2}\otimes\mathbf{C}_{\lambda}.$
Similarly, for $\mu\in\mathbf{C}$ there is a unique $\widehat{\mathsf{E}}_{Y}$-module $\mathsf{M}^{\mu}_{M}$ simple along $\Lambda_{M}$ such that
$i_{M}^{*}\mu_{\Lambda_{M}}^{dR}(\mathsf{M}^{\mu}_{M})\cong K_{M}^{1/2}\otimes\mathbf{C}_{\mu}.$
The goal of this section is to prove the following result.
@@ 7.1. Statement
$\mathsf{M}^{\lambda}_{L}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{S}}\mathsf{M}^{\mu}_{M}\cong\mathrm{RH}^{-1}(\phi_{L\cap M}\otimes\mathbf{C}_{\mu+\lambda}),$
@@
Corollary 7.2.
Let $\mu\in\mathbf{C}$. There is an isomorphism
$i_{L}^{*}\mu^{dR}_{\Lambda_{L}}\mathsf{M}^{\mu}_{M}\cong\phi^{\mathcal{D}}_{L\cap M}\otimes\mathbf{C}_{\mu}$
of $\mathcal{D}_{L}$-modules equipped with automorphisms, where we identify $\mathcal{D}^{\sqrt{v}}_{L}$-modules with $\mathcal{D}_{L}$-modules using the orientation data $K_{L}^{1/2}$.
Proof.
Pick any $\lambda\in\mathbf{C}$. By 7.1 we have
$\mathsf{M}^{\lambda}_{L}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{S}}\mathsf{M}^{\mu}_{M}\cong\mathrm{RH}^{-1}(\phi_{L\cap M}\otimes\mathbf{C}_{\mu+\lambda}).$
By 4.13 we have
$\mathsf{M}^{\lambda}_{L}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{S}}\mathsf{M}^{\mu}_{M}\cong\mathrm{RH}^{-1}\left((\mu^{dR}_{\Lambda_{L}}\mathsf{M}^{\lambda}_{L}\otimes^{\mathbb{L}}_{\mathcal{D}^{\sqrt{v}}_{\tilde{\Lambda}_{L}}}\mu^{dR}_{\Lambda_{L}}\mathsf{M}^{\mu}_{M})|_{L}[-1]\right).$
By definition $i^{*}\mu^{dR}_{\Lambda_{L}}\mathsf{M}^{\lambda}_{L}\cong K_{L}^{1/2}$ equipped with the automorphism $\exp(2\pi i\lambda)$. Therefore, we obtain an isomorphism
$\mathbf{C}_{\mu+\lambda}\otimes\phi_{L\cap M}\cong\mathbf{C}_{\lambda}\otimes(K_{L}\otimes^{\mathbb{L}}_{\mathcal{D}_{\tilde{\Lambda}_{L}}}\mu^{dR}_{\Lambda_{L}}\mathsf{M}^{\mu}_{M})|_{L}[-1]$
of perverse sheaves on $L$ equipped with an automorphism. Identifying $D$-modules with perverse sheaves via the Riemann–Hilbert correspondence we get the result. ∎
Throughout this section we let $\dim L=\dim M=n$.
@@
The manifold $S$ admits local polarizations. We use them to get a local description of $(Y\rightarrow S,\Lambda_{L}\rightarrow L,\Lambda_{M}\rightarrow M)$ following the terminology from [Bussi, Section 2]:
@@ 7.3. Local description of DQ modules
  • In the $L$-chart we have $\widehat{\mathsf{E}}_{Y}=\widehat{\mathcal{E}}^{\sqrt{v}}_{L\times\mathbf{C}}$ and $\widehat{\mathsf{W}}_{S}=\widehat{\mathcal{W}}^{\sqrt{v}}_{L}$. The modules are
@@
  • In the $M$-chart we have $\widehat{\mathsf{E}}_{Y}=\widehat{\mathcal{E}}^{\sqrt{v}}_{M\times\mathbf{C}}$ and $\widehat{\mathsf{W}}_{S}=\widehat{\mathcal{W}}^{\sqrt{v}}_{M}$. The modules are
@@
We use the gluing data for the perverse sheaf $\phi_{L\cap M}$ as described in [Bussi, Section 2.3]. Cover $S$ by open sets which admit polarizations and consider them as $L$-charts. Given an intersection of two such $L$-charts $U_{1},U_{2}$, we may cover $U_{1}\cap U_{2}$ by open sets $\{V_{a}\}$ with polarizations such that for every $a$ the polarization on $V_{a}$ is transverse both to the polarization on $U_{1}$ and on $U_{2}$. We regard $V_{a}$ as $M$-charts.
@@
It is enough to check commutativity of the diagram locally. Consider a point $x\in L\cap M$. We have that $\Xi\colon(L\cap M,L,f,i)\rightarrow(L\cap M,L\times M,h,i)$ is a critical embedding. Therefore, by [JoycedCrit, Proposition 2.22] in a neighborhood of $x$ there is a local isomorphism $L\times M\cong L\times W$, where $W=\mathrm{T}_{x}M$ is equipped with the Maslov quadratic form $q=q(\mathrm{T}_{M},\mathrm{T}_{\pi_{1}},\mathrm{T}_{\pi_{2}})\colon W\rightarrow\mathbf{C}$, such that under this isomorphism $h\mapsto f\boxplus q$ and the embedding $\Xi\colon L\rightarrow L\times M$ becomes $(\mathrm{id}\times 0)\colon L\rightarrow L\times W$. The $\mathbf{Z}/2$-torsor $P_{\Xi}$ in a neighborhood of $x$ canonically identifies with $P_{q}$.
@@ 7.5. Example: a clean Lagrangian intersection
described in [JoycedCrit, Theorem 2.28 (i)] for a general d-critical locus. By [Bussi, Theorem 3.1] we may additionally identify